# zkLoRA: Fine-Tuning Large Language Models with Verifiable Security via Zero-Knowledge Proofs

### v1

@@ III-A1 Transformer Layers
Each Transformer layer consists of two main components: a multi-head attention (MHA) sub-layer [28] and a feed-forward multilayer perceptron (MLP) sub-layer. Let $x\in\mathbb{R}^{n}$ denote the input token sequence to the LLM, and $X_{\ell}\in\mathbb{R}^{n\times d}$, $\ell=1,\cdots,L$, denote the output of the $\ell^{\text{th}}$ Transformer layer, where $n$ represents the sequence length and $d$ represents the hidden dimension. Since the $L$ Transformer layers are sequentially cascaded, $X_{\ell-1}$ serves as the input to the $\ell^{\text{th}}$ Transformer layer for $\ell=2,\cdots,L$. The input token sequence $x$ is the input to the first Transformer layer. Below, we describe the structure and computations of the $\ell^{\text{th}}$ Transformer layer, which takes $X_{\ell-1}$ as input and computes $X_{\ell}$ as output.
@@
Before the MLP sub-layer, LayerNorm is applied to normalize the residual-enhanced representation $R_{\ell}=X_{\ell-1}+O_{\ell}$: $R^{\prime}_{\ell}=\mathsf{LayerNorm}(R_{\ell})$. The output from LayerNorm $R^{{}^{\prime}}_{\ell}$ is then passed through the feed-forward MLP with SwiGLU activation [2]:
$\begin{array}[]{l}\mathsf{FFN}(R^{\prime}_{\ell})=\left(\phi(R^{\prime}_{\ell}W_{{gate},\ell}^{\top})\odot\right.\left.R^{\prime}_{\ell}W_{{up},\ell}^{\top}\right)W_{{down},\ell}^{\top}\end{array}$ (12)
@@ III-B LoRA Fine Tuning
Instead of directly updating the parameter matrices as in standard fine-tuning, LoRA [3] offers a parameter-efficient alternative by freezing the original parameter matrices and introducing trainable low-rank matrices into the architecture. These low-rank matrices are added to the original parameter matrices during both the forward and backward propagation phases of LLM fine-tuning. Mathematically, for a target matrix $W_{\ell}\in\{W_{Q,\ell},W_{K,\ell},W_{V,\ell},W_{{up},\ell},W_{{down},\ell},W_{{gate},\ell}\}$, LoRA modifies it in each iteration as: $W^{\prime}_{\ell}=W_{\ell}+\Delta W_{\ell}$ with $\Delta W_{\ell}=B_{\ell}A_{\ell}$. Here, $\Delta W_{\ell}$ represents the fine-tuning update, which is decomposed into two small low-rank matrices $B_{\ell}\in\mathbb{R}^{d\times r}$ and $A_{\ell}\in\mathbb{R}^{r\times k}$, with rank $r$ satisfying $r\ll\min{\left(d,k\right)}$. Since LoRA fine-tuning only the low-rank matrices $B_{\ell}$ and $A_{\ell}$, while keeping the original parameter matrices $W_{\ell}$ frozen, the number of trainable parameters is significantly reduced compared to standard fine-tuning. This makes LoRA an efficient and scalable method for adapting large language models.
@@ III-C Gradient Computations
In the following sections, we will demonstrate how our proposed zkLoRA achieves verifiable computations throughout the entire LoRA procedure using zero-knowledge proofs.
IV Design of zkLoRA
The goal of zkLoRA is to introduce zero-knowledge verifiability into LoRA fine-tuning for LLMs. In zkLoRA, a prover $\mathcal{P}$ is responsible for performing the computation required for LoRA fine-tuning on the LLM and generating cryptographic proofs that ensure all computational steps strictly adhere to the predefined procedures without deviation. Meanwhile, a verifier $\mathcal{V}$ can efficiently validate these proofs without accessing any privacy-sensitive information, such as model parameters or training data. Owing to the zero-knowledge nature of the proofs, the verifier can confirm whether the LoRA fine-tuning process was executed correctly, without extracting any additional sensitive details.
@@ IV Design of zkLoRA
As discussed in Section III-B, the entire computation procedure of LoRA can be divided into three sequential phases, and zkLoRA follows the same structure. In Section IV-A, we detail the key technical aspects of the backward propagation phase in zkLoRA as a representative example to illustrate how non-arithmetic operations are proven. The proving scheme for non-arithmetic operations in the forward propagation and parameter update phases follows the same approach. In Section IV-B, we will put everything together to present the complete zkLoRA framework.
IV-A zkLoRA for Backward Propagation Phase
During the backward propagation phase, the gradients of the cross-entropy loss $\mathcal{L}^{t}$ wrt the parameters are computed and recursively propagated backward through each Transformer layer $\ell$, starting from the final output layer. For the backward propagation phase at each Transformer layer, we essentially aim to prove the gradient computation process expressed by (LABEL:eq:grad_ol)-(31). The arithmetic operations in the backward propagation phase such as matrix additions and multiplications, are the same as the forward propagation phase, and we can encode the relevant matrices as MLEs and then prove the computation correctness using the sumcheck protocol, as proposed in [29, 7]. In contrast to the forward propagation phase, we face increased complexity in the backward propagation phase due to the gradient computations, which involve more non-arithmetic functions.
@@
We first establish the zkSwiGLU protocol, which achieves a zero-knowledge proof for the computations of the SwiGLU activation $\phi(\cdot)$ and its derivative $\phi^{\prime}(\cdot)$. Note that both the SwiGLU activation $\phi(\cdot)$ and its derivative $\phi^{\prime}(\cdot)$ apply a non-linear mapping to their input matrices in an element-wise manner. Since the proof protocols for $\phi(\cdot)$ and $\phi^{\prime}(\cdot)$ are identical in zkLoRA, we present the details of the proof protocol for $\phi(\cdot)$ as a representative.
@@
    Activation Correctness: The secret set $S$ is designated as $S=\{z^{\prime}_{i}+\alpha g_{i}\}_{i=1}^{D}$, where the random linear combination coefficient $\alpha\in\mathbb{F}$ is chosen by the verifier  [^1] To ensure the protocol is non-interactive, zkLoRA leverages the Fiat-Shamir heuristic, which eliminates the need for direct interaction between the prover and verifier. Specifically, each verifier challenge is computed by hashing all previously generated commitments and public messages in the protocol, making the protocol non-interactive.. The predefined set $T$ is assigned as $T=\{x+\alpha\phi(x)\mid x\in\mathbb{F}_{B}\}$. Here, in $T$, each $x$ represents a possible value of the quantized activation input $z^{\prime}_{i}$, and $\phi(x)$ denotes the corresponding activation output. The satisfaction of this lookup argument ensures the correctness of the computation relationship between the activation function’s input and output, rather than any other arbitrary computations.
@@
    Quantization Correctness: The secret set $S$ is designated as $S=\{\zeta z^{\prime}_{i}+r_{i}\}_{i=1}^{D}$, while the predefined set $T$ is assigned as $T=\{\zeta x+r\mid x\in\mathbb{F}_{B},r\in[-\zeta/2,\zeta/2)\}$. Here, in $T$, $x\in\mathbb{F}_{B}$ corresponds to a possible value of the quantized input element to the activation function (i.e., the possible value of $z^{\prime}_{i}$), and $r\in[-\zeta/2,\zeta/2)$ corresponds to a possible value of the quantization residue $r_{i}$. Here, in $T$, $x\in\mathbb{F}_{B}$ corresponds to a possible value of the quantized input element to the activation function (i.e., the possible value of $z^{\prime}_{i}$), and $r\in[-\zeta/2,\zeta/2)$ represents the quantization residue. The satisfaction of this lookup argument ensures the correctness of the quantization process from the original value $z_{i}$ to its quantized value $z^{\prime}_{i}$, for $i\in[D]$.
@@
Protocol 2 zkLoRA: Zero-Knowledge Proof Framework for Verifiable LoRA Fine-tuning on LLMs
@@
We formally present the complete zkLoRA protocol, a zero-knowledge proof framework designed specifically to enable verifiable LoRA fine-tuning on Transformer based LLMs. zkLoRA rigorously proves correctness at each computational step of LoRA fine-tuning, encompassing the forward and backward propagations through the Transformer layers, as well as the parameter update.
@@
Protocol 2 presents zkLoRA as a pseudo code, integrating the above protocols into a complete proving procedure for LoRA fine turning.
@@ V Security Analysis
In this section, we present a formal security analysis of the zkLoRA protocol, focusing on its completeness, soundness, and zero-knowledge properties under standard cryptographic assumptions.
The zkLoRA protocol is designed and implemented using the Hyrax polynomial commitment scheme, whose security guarantees are thoroughly discussed in [30]. To achieve non-interactive proof functionality, the protocol leverages the Fiat–Shamir heuristic, which operates under the standard model of a random oracle. This approach aligns with the formal frameworks commonly adopted in non-interactive proof systems [30]. Additionally, we assume the use of a sufficiently large finite field $\mathbb{F}$, typically with $|\mathbb{F}|\approx 2^{254}$, such as that derived from the BLS12-381 elliptic curve.
@@ V-A Soundness
Statement: zkLoRA ensures soundness, meaning that a verifier will accept proofs for incorrectly computed operations from a malicious PPT prover who deviates from the zkLoRA protocol in Protocol 2 only with a negligible probability of $\varepsilon_{\mathrm{sound}}\leq\frac{m\cdot d_{\mathrm{max}}}{|\mathbb{F}|}+\frac{C}{|\mathbb{F}|}+\varepsilon_{\mathrm{binding}}$, where $m$ is the number of iterations rounds in the sumcheck protocol, $d_{max}$ is the largest degree with respect to any single variable among all polynomials involved in the sumcheck protocol, $C$ is the number of elements checked by a lookup argument in a non-arithmetic operation (e.g., element-wise products, the SwiGLU activations and its gradient computation, the Softmax operations, and the matrix transpositions), $\varepsilon_{\mathrm{binding}}$ is the binding error probability for Hyrax-style polynomial commitments.
Analysis: We analyze the soundness errors of zkLoRA in its arithmetic computations, non-arithmetic computations, and polynomial commitments, respectively.
@@
  • Arithmetic Operations: zkLoRA proves the arithmetic operations in LoRA (specifically, the matrix multiplications and additions) by representing them as MLEs and leveraging the sumcheck protocol for proof generation. According to the Schwartz-Zippel lemma and the analysis result in [26], a false summation result over the polynomial is accepted with probability at most $m\cdot d_{\mathrm{max}}/|\mathbb{F}|$.
@@
  • Non-arithmetic Operations: zkLoRA proves non-arithmetic computations in LoRA (e.g., element-wise products, SwiGLU activations, Softmax operations, and matrix transposes) using lookup-based arguments and sumcheck protocols. The prover employs the sumcheck protocol to verify rational-function identities, ensuring set-membership conditions ($S\subseteq T$) are satisfied. Specifically, the sum $\sum_{i=1}^{D}\frac{1}{X+s_{i}}$ equals $\sum_{j=1}^{N}\frac{m_{j}}{X+t_{j}}$, where $s_{i}$ are elements of the secret tensor $S$, $t_{j}$ are elements of the lookup table $T$, $m_{j}$ is the multiplicity of $t_{j}$ in $S$, and $X$ is a random challenge from $\mathbb{F}$. Any mismatch is detected with probability at most $C/|\mathbb{F}|$, ensuring negligible success for malicious provers.
@@
  • Polynomial Commitments: The Hyrax-style polynomial commitments utilized in zkLoRA are computationally binding, ensuring that once a polynomial is committed, the prover cannot open it to a different value except with negligible probability $\varepsilon_{\mathrm{binding}}$. A violation of the binding property would constitute a breach of the discrete logarithm assumption in the random oracle model [14], which is considered infeasible in practice.
@@
In summary, zkLoRA ensures that the soundness error for any invalid computation is bounded by the sum of the three components described above. As a result, the overall soundness error remains negligible, guaranteeing that a malicious prover has an exceptionally low probability of persuading the verifier to accept an incorrect computation.

### v3

@@ III-A1 Transformer Layers
Each Transformer layer consists of two main components: a multi-head attention (MHA) sub-layer [vaswani2017attention] and a feed-forward multilayer perceptron (MLP) sub-layer. Let $x\in\mathbb{R}^{n}$ denote the input token sequence to the LLM, and $X_{\ell}\in\mathbb{R}^{n\times d}$, $\ell=1,\cdots,L$, denote the output of the $\ell^{\text{th}}$ Transformer layer, where $n$ represents the sequence length and $d$ represents the hidden dimension. Since the $L$ Transformer layers are sequentially cascaded, $X_{\ell-1}$ serves as the input to the $\ell^{\text{th}}$ Transformer layer for $\ell=2,\cdots,L$. The input token sequence $x$ is the input to the first Transformer layer. Below, we describe the structure and computations of the $\ell^{\text{th}}$ Transformer layer, which takes $X_{\ell-1}$ as input and computes $X_{\ell}$ as output.
@@
Before the MLP sub-layer, LayerNorm is applied to normalize the residual-enhanced representation $R_{\ell}=X_{\ell-1}+O_{\ell}$: $R^{\prime}_{\ell}=\mathsf{LayerNorm}(R_{\ell})$. The output from LayerNorm $R^{{}^{\prime}}_{\ell}$ is then passed through the feed-forward MLP with SwiGLU activation [touvron2023llama]:
$\ignorespaces\mathsf{FFN}(R^{\prime}_{\ell})=\left(\phi(R^{\prime}_{\ell}W_{{gate},\ell}^{\top})\odot\right.\left.R^{\prime}_{\ell}W_{{up},\ell}^{\top}\right)W_{{down},\ell}^{\top}$ (12)
@@ III-B LoRA Fine Tuning
Instead of directly updating the parameter matrices as in standard fine-tuning, LoRA [hu2022lora] offers a parameter-efficient alternative by freezing the original parameter matrices and introducing trainable low-rank matrices into the architecture. These low-rank matrices are added to the original parameter matrices during both the forward and backward propagation phases of LLM fine-tuning. Mathematically, for a target matrix $W_{\ell}\in\{W_{Q,\ell},W_{K,\ell},W_{V,\ell},W_{{up},\ell},W_{{down},\ell},W_{{gate},\ell}\}$, LoRA modifies it in each iteration as: $W^{\prime}_{\ell}=W_{\ell}+\Delta W_{\ell}$ with $\Delta W_{\ell}=B_{\ell}A_{\ell}$. Here, $\Delta W_{\ell}$ represents the fine-tuning update, which is decomposed into two small low-rank matrices $B_{\ell}\in\mathbb{R}^{d\times r}$ and $A_{\ell}\in\mathbb{R}^{r\times k}$, with rank $r$ satisfying $r\ll\min{\left(d,k\right)}$. Since LoRA fine-tuning only the low-rank matrices $B_{\ell}$ and $A_{\ell}$, while keeping the original parameter matrices $W_{\ell}$ frozen, the number of trainable parameters is significantly reduced compared to standard fine-tuning. This makes LoRA an efficient and scalable method for adapting large language models.
@@ III-C Gradient Computations
In the following sections, we will demonstrate how our proposed VeriLoRA achieves verifiable computations throughout the entire LoRA procedure using zero-knowledge proofs.
IV Design of VeriLoRA
The goal of VeriLoRA is to introduce zero-knowledge verifiability into LoRA fine-tuning for LLMs. In VeriLoRA, a prover is responsible for performing the computation required for LoRA fine-tuning on the LLM and generating cryptographic proofs that ensure all computational steps strictly adhere to the predefined procedures without deviation. Meanwhile, a verifier can efficiently validate these proofs without accessing any privacy-sensitive information, such as model parameters or training data. Owing to the zero-knowledge nature of the proofs, the verifier can confirm whether the LoRA fine-tuning process was executed correctly, without extracting any additional sensitive details.
@@ IV Design of VeriLoRA
As discussed in Section III-B, the entire computation procedure of LoRA can be divided into three sequential phases, and VeriLoRA follows the same structure. In Section IV-A, we detail the key technical aspects of the backward propagation phase in VeriLoRA as a representative example to illustrate how non-arithmetic operations are proven. The proving scheme for non-arithmetic operations in the forward propagation and parameter update phases follows the same approach. In Section IV-B, we will put everything together to present the complete VeriLoRA framework.
IV-A VeriLoRA for Backward Propagation Phase
During the backward propagation phase, the gradients of the cross-entropy loss $\mathcal{L}^{t}$ wrt the parameters are computed and recursively propagated backward through each Transformer layer $\ell$, starting from the final output layer. For the backward propagation phase at each Transformer layer, we essentially aim to prove the gradient computation process expressed by (LABEL:eq:grad_ol)-(31). The arithmetic operations in the backward propagation phase such as matrix additions and multiplications, are the same as the forward propagation phase, and we can encode the relevant matrices as MLEs and then prove the computation correctness using the sumcheck protocol, as proposed in [sun2024zkdl, sun2024zkllm]. In contrast to the forward propagation phase, we face increased complexity in the backward propagation phase due to the gradient computations, which involve more non-arithmetic functions.
@@ IV-A VeriLoRA for Backward Propagation Phase
We first establish the zkSwiGLU protocol, which achieves a zero-knowledge proof for the computations of the SwiGLU activation $\phi(\cdot)$ and its derivative $\phi^{\prime}(\cdot)$. Note that both the SwiGLU activation $\phi(\cdot)$ and its derivative $\phi^{\prime}(\cdot)$ apply a non-linear mapping to their input matrices in an element-wise manner. Since the proof protocols for $\phi(\cdot)$ and $\phi^{\prime}(\cdot)$ are identical in VeriLoRA, we present the details of the proof protocol for $\phi(\cdot)$ as a representative.
@@
    Activation Correctness: The secret set $S$ is designated as $S=\{z^{\prime}_{i}+\alpha g_{i}\}_{i=1}^{D}$, where the random linear combination coefficient $\alpha\in\mathbb{F}$ is chosen by the verifier  [^1] To ensure the protocol is non-interactive, modern cryptographic protocols of this type typically leverage the Fiat-Shamir heuristic, which eliminates the need for direct interaction between the prover and verifier. Specifically, each verifier challenge is computed by hashing all previously generated commitments and public messages in the protocol, making the protocol non-interactive.. The predefined set $T$ is assigned as $T=\{x+\alpha\phi(x)\mid x\in\mathbb{F}_{B}\}$. Here, in $T$, each $x$ represents a possible value of the quantized activation input $z^{\prime}_{i}$, and $\phi(x)$ denotes the corresponding activation output. The satisfaction of this lookup argument ensures the correctness of the computation relationship between the activation function’s input and output, rather than any other arbitrary computations.
@@
    Quantization Correctness: The secret set $S$ is designated as $S=\{\zeta z^{\prime}_{i}+r_{i}\}_{i=1}^{D}$, while the predefined set $T$ is assigned as $T=\{\zeta x+r\mid x\in\mathbb{F}_{B},r\in[-\zeta/2,\zeta/2)\}$. Here, in $T$, $x\in\mathbb{F}_{B}$ corresponds to a possible value of the quantized input element to the activation function (i.e., the possible value of $z^{\prime}_{i}$), and $r\in[-\zeta/2,\zeta/2)$ corresponds to a possible value of the quantization residue $r_{i}$. The satisfaction of this lookup argument ensures the correctness of the quantization process from the original value $z_{i}$ to its quantized value $z^{\prime}_{i}$, for $i\in[D]$.
@@
Protocol 2 VeriLoRA: Zero-Knowledge Proof Framework for Verifiable LoRA Fine-tuning on LLMs
@@
We formally present the complete VeriLoRA protocol, a zero-knowledge proof framework designed specifically to enable verifiable LoRA fine-tuning on Transformer based LLMs. VeriLoRA rigorously proves correctness at each computational step of LoRA fine-tuning, encompassing the forward and backward propagations through the Transformer layers, as well as the parameter update.
@@
To integrate these protocols for proving different operations of the VeriLoRA fine tuning on LLM, we follow the same sequential-composition strategy as zkLLM [sun2024zkllm]: i) Fixing random point-value claims through sumcheck/lookup, thereby binding the prover to specific evaluation points and claimed values; ii) Requiring each subsequent proving step to justify exactly the same points via further randomized reductions (e.g., random aggregation), conditioned on success in prior protocols; iii) Finalizing by opening evaluations against the same committed parameters, leveraging the binding property of the commitment scheme. These steps prevent the prover from reordering the proving steps or retroactively altering intermediate values, as both challenges and commitments are binding and fresh across proving steps.
Protocol 2 presents VeriLoRA as a pseudo code, integrating the above protocols into a complete proving procedure for LoRA fine turning; subproofs are assembled in the arithmetic circuit’s reverse logical order.
@@ V Security Analysis
In this section, we present a formal security analysis of the VeriLoRA protocol, focusing on its completeness, soundness, and zero-knowledge properties under standard cryptographic assumptions.
The VeriLoRA protocol is designed and implemented using the Hyrax polynomial commitment scheme, whose security guarantees are thoroughly discussed in [raffel2020exploring]. To achieve non-interactive proof functionality, the protocol leverages the Fiat–Shamir heuristic, which operates under the standard model of a random oracle. This approach aligns with the formal frameworks commonly adopted in non-interactive proof systems [raffel2020exploring]. Additionally, we assume the use of a sufficiently large finite field $\mathbb{F}$, typically with $|\mathbb{F}|\approx 2^{254}$, such as that derived from the BLS12-381 elliptic curve.
@@ V-A Soundness
Statement: VeriLoRA ensures soundness, meaning that a verifier will accept proofs for incorrectly computed operations from a malicious PPT prover who deviates from the computations specified in Protocol 2 only with negligible probability. This probability is bounded (using a union bound with conditional probabilities) by: $\varepsilon_{\mathrm{sound}}\leq\sum_{j}\varepsilon_{j}$, where $\varepsilon_{j}$ represents the soundness error of each proving step, conditioned on the correctness of all prior steps. Each $\varepsilon_{j}$ is defined by the following values: $\frac{md_{\mathrm{max}}}{|\mathbb{F}|}$ (for arithmetic operations), $\frac{C}{|\mathbb{F}|}$ (for non-arithmetic operations), and $\varepsilon_{\mathrm{binding}}$ (for polynomial commitments). Here, $m$ denotes the number of iteration rounds in the sumcheck protocol, $d_{\mathrm{max}}$ is the largest degree with respect to any single variable among all polynomials involved in the sumcheck protocol, $C$ represents the number of elements verified by a lookup argument in a non-arithmetic operation (e.g., element-wise products, SwiGLU activations and their gradient computations, Softmax operations, and matrix transpositions), and $\varepsilon_{\mathrm{binding}}$ refers to the binding error probability of Hyrax-style polynomial commitments.
Analysis: The soundness property of VeriLoRA is upheld by the protocol’s sequential composition strategy that prevents the prover from reordering proving steps or retroactively modifying intermediate values, as both challenges and commitments are binding and freshly generated at each proving step. Consequently, the overall soundness error of VeriLoRA is bounded by the union bound with conditional probabilities (as given above in the statement), which represents the standard upper bound for sequential composition without assuming independence between proving steps. We analyze each $\varepsilon_{j}$, which represents the soundness error of VeriLoRA at each proving step, for arithmetic computations, non-arithmetic computations, and polynomial commitments, respectively.
@@
  • Arithmetic Operations: VeriLoRA proves the arithmetic operations in LoRA (specifically, the matrix multiplications and additions) by representing them as MLEs and leveraging the sumcheck protocol for proof generation. According to the Schwartz-Zippel lemma and the analysis result in [thaler2022proofs], a false summation result over the polynomial is accepted with probability at most $md_{\mathrm{max}}/|\mathbb{F}|$.
@@
  • Non-arithmetic Operations: VeriLoRA proves non-arithmetic computations in LoRA (e.g., element-wise products, SwiGLU activations, Softmax operations, and matrix transposes) using lookup-based arguments and sumcheck protocols. The prover employs the sumcheck protocol to verify rational-function identities, ensuring set-membership conditions ($S\subseteq T$) are satisfied. Specifically, the sum $\sum_{i=1}^{D}\frac{1}{X+s_{i}}$ equals $\sum_{j=1}^{N}\frac{m_{j}}{X+t_{j}}$, where $s_{i}$ are elements of the secret tensor $S$, $t_{j}$ are elements of the lookup table $T$, $m_{j}$ is the multiplicity of $t_{j}$ in $S$, and $X$ is a random challenge from $\mathbb{F}$. Any mismatch is detected with probability at most $C/|\mathbb{F}|$, ensuring negligible success for malicious provers.
@@
  • Polynomial Commitments: The Hyrax-style polynomial commitments utilized in VeriLoRA are computationally binding, ensuring that once a polynomial is committed, the prover cannot open it to a different value except with negligible probability $\varepsilon_{\mathrm{binding}}$. A violation of the binding property would constitute a breach of the discrete logarithm assumption in the random oracle model [wahby2018doubly], which is considered infeasible in practice.
@@
In summary, VeriLoRA ensures that the soundness error for any invalid computation is bounded by the sum of the these error components described above. As a result, the overall soundness error remains negligible, guaranteeing that a malicious prover has an exceptionally low probability of persuading the verifier to accept an incorrect computation.